OpenStudy (anonymous):

How does one solve sin(2x)=sqrt(3)/2? Basically, I am unsure how to approach a multiple within the sin definition?

OpenStudy (anonymous):

use sinx = sina then x= n(pi) + )(-1)^n (a) wher n is any integer...

OpenStudy (zzr0ck3r):

so \[2x=sin^{-1}(\frac{\sqrt{3}}{2})\]

OpenStudy (anonymous):